Tom is 1/3 Joe's age. David is 1/2 Joe's age. Who is the youngest?
Answer by stanbon(75887)

Tom is 1/3 Joe's age. David is 1/2 Joe's age. Who is the youngest?
----
t = (1/3)j
d = (1/2)j
-----------------
1/2 - 1/3 = (3-2)/6 = 1/6
Since the difference is positive, 1/2 is larger than 1/3.
Ans: Tom is the youngest.
Cheers,
Stan H.
============